Examples on Elements of Christian and Pagan Beliefs in Irish Names of Places

GA4116 Essay

Oliver Gerler

9954465 GA4116: Irish Language 2

University of Limerick, Ireland

May 11, 2000

The question of the roots of names of places (not only) in Ireland is a difficult and interesting one. Interesting because of the reflections of history that can be seen in the change of particular names over time. According to [4] the changings and roots of the names can be divided in several categories. One of it, the systematic changes, are quite traceable, as the old and original Irish names are either still in use^{1} or can still be heard out of (though sometimes strange) transcriptions into the english language.

Difficult because of the many tribes and peoples time has swept to the shores of this island and beyond. Over centuries, Norsemen, Danes, Frenchmen and not to mention the inhabitants of the next island to the east have changed the look and feel of Ireland. According to [3] (and as shown by some examples below) many a names are not from Gaelic roots but are indeed due to the invaders (but quite often not correctly retranslated into english or Irish which adds up to the difficulties in tracing their original meanings).

Nonetheless, the Gaelic roots persists and they reflect not only in the still used old language but in the place-names, too. But the greatest influence on the forming of the placenames have the old pagan myths and beliefs and the newer Christian influences, started by St.Patrick (or what is made of him by projecting several different missionaries onto one person, as research has brought up of what could have happened).

There can be found many names of places that reflect connections to Christian beliefs, but still there are a lot of the old names continuing to exist. These old names can tell us about the people who lived there and how they lived there. They can tell us a lot about their beliefs and traditions and not so much about their daily life (which becomes obvious, as daily life is not that interesting to record and to pass on through the generations than special events^{2}).

Placenames^{3} in Ireland are not only derived from religious backgrounds, but more often than not just simply from geographically situations and sites at that place. Two examples may elaborate that:

Phoenix Park: this park in Dublin near the Viceregal Lodge took its name from a spring called Fionn uisg’, clear or limpid water.

Carlow: this town is called Cetherloch in old Irish documents, which means quadruple lake as the barrow formed four lakes there in old times.

Other examples are Inishmore (Inis Móre) ’large island’ or just Knock (cnoc) ’hillock’.

1[6] states the law from 1962 that place-names shall be shown in Irish and English or at least in Irish only. It also states that in order to find the original Irish names local experts and societies are required and to be consulted.

2Of that arises the question on how accurate archaeologists can be in reconstructing everyday routines. Many peoples occur like very religious and ritualistic groups, but that is only one side of their world.

3All of the citations and explanations have either been taken from [3] or [4] and will not be referenced furthermore. The transcriptions itself follow the listings in these references and are not changed to modern Irish.

Another major influence on the names came from the invasions from the North-East by the Danes and Norsemen. Many of these descriptive names have been mistranslated in later times, prominent examples of this may be the following:

Dalkey Island: The Irish name is Delg inis ’thorn island’, which the Danes who had a fortress on it, translated it to their own Dalk Ey with the same meaning.

Ireland’s Eye: The same happened here by taking the sound of the Danish ey and anglizising it by taking Eye. Furthermore the name has nothing to do with Ireland, as its original name Inis Ereann plainly means the island of Eria, a woman’s name.

Leinster, Munster, Ulster: The termination ster (staðr), Scandinavian for ’place’, is added to Laighen, Mumhan and Uladh to form the actual names of these three provinces.

Christian elements in Irish place names are easily found. The syllable ’kill’ for example has its root in cill, ’church’. According to [3, p. 67] the name Kilmurry (Cill Mhuire), ’the church of Mary’, can be found more than 50 times throughout Ireland. Other examples are Killbreedy (Cill Bhrighde) ’church of St. Brigid’ or Shankill (Seán Cill) ’old church’. In Killadysart the second part is derived from disert ’hermitage’ which occurs as Desert, Dysart or sometimes in the forms Dysert, Isert or even Ishart. Other examples may be Killedy or Kilmeedy, both meaning ’church of Ide’ or Kilkenny (Cill Chainnigh) ’church of St. Canice’; but Killery has another origin, corrupted by a change of l to r, Cael shaile ruadh, meaning a ’reddish sea inlet’.

Other part of names like Teampall or the anglicized Temple can be found in various names like Templemore (An Teampall Mór) ’big church’. From the Greek (and therefore Latin) ϵκλϵσια stem comes the syllable eaglais, in forms like Aglish, Eglish or even Heagles - all meaning ’church’. Baisleac, meaning ’basilica’, occurs in the name Baslick, resembling the correct pronounciation. Donagh or Réilig is used for church, too, as in Donaghedy (Domhnach Chaeide) ’St.Caida’s church’ or Donnybrook (Domhnach Broc) ’St. Broc’s church’. The Irish word Donagh from Latin dominicus means ’belonging to the Lord’ and corresponds to the Greek κιριακων, the ultimate origin of ’church’. In the Inverness-Shire in Scotland places like Reelick or Ruilick can be found, related to the Irish réilig and in Wales places containing Llan, like Llandudno ’church of Tudno’ or Llanfair ’church of Mary’, are very common also^{4}.

In many names of places the part ’abbey’ can be found, like Abbeyfeale (Mainistir na Feile), this time a correct translation to the english word, or Abbeylara (Mainistir leath rath).

4It may be interesting to note that according to [5] St. Bridget is the only saint who receives the title of saint in Welsh placenames as in Llansantffraed ’church of St. Bride’.

In many forms english transcriptions for baile ’town’ can be found, like Balbriggan (Baile Breacain) ’town of St. Brecan’ or Ballynakill (Baile na Cill) ’town of the church’, meaning the main town of the area, similar to Ballynahaglish with the same meaning. Ballysagart therefore is the ’town of the priest’.

In the case of Caldragh (Cealdrach) ’burial ground’ it can not be said without further investigation in the area itself if this particular name comes from a Christian grave yard or if there already has been one from the pagan times.

Croaghpatrick ’St. Patrick’s hill’ is one example of many places claiming to have had connections to the old saint. But it is not sure either if this was one man alone or a group of wandering preachers and monks, and so the claims can be justified or not.

As there are a lot of different saints who all have their several placenames accounted to themselves, not only the important ones got churches, hills, wells, hollows and other geographically places named after them, but also lower saints and other "important" people as for example Liadhan, who is believed to be Ciarán’s mother, in Cill Liadhan or Adhamhman, who wrote "The Life of St. Colm Cille", in Toberawnawn ’the well of Adhamhnam’ or Aireceal Adhamhnam.

Pagan associations in place names are often to be found, too, despite the longer time that has passed since these times. They have a closer connection to nature than to religious topics. Especially some important festivals and the (one) place of their happening are recorded, like for Bealtaine or Tailtenn. Other festivals like Samhain or Lunasa are of relevance to the lifes of the pagans, too. The place of their dwelling often occurs in the names of the places and of the different tribes that claimed the areas their own. Some examples for that kind of place-names:

Beltany (Bealtaine) ’May’, especially the beginning of May. It is believed that king Uisneach Tuathaill has started this festival in honour of tene-bil (old-Irish for ’fire-god’ or the ’fire of the gods’), who was said to prevent diseases in cattle. Lisbalting ’The fort of Bealtaine’ is one place where the name exists.

The big festival of Lunasa (Lugh nasa) ’gathering of Lui’ is said to be started by Lui King of the Tuatha De Danann. See Cooley Hills and Teltown below.

One of the most widespread festival (from Europe to America and beyond) will be Samhain ’Day of the Dead’ (but literally often explained as samh fhuin ’end of summer’) where the end of summer will be celebrated on the 1st of November. The center of the festivities (as it was also instituted by Tuathaill at Tlachtga in Meath) is the feis at the seat of the highkings in Tara. According to lure and legend the festival goes on for 3 days and nights before and after that day where fires are lit and games and sports are carried on. Several myths are built around that festival that is now minimized as a play for children in the day of Halloween. Some places containing that name are Knocksouna (Cnoc Samhna) ’hill of Samhain’ or Mullasawny (sometimes -sunny) (Mullach Samhna) ’hill-summit of Samhuin’ and Drumhawan (Druim Shamhauin) ’the ridge of Samhuin’.

Castledermot (Disert Diarmaid) ’hermitage of Diarmaid’. Diarmaid is well known from the tale Tóruigheacht Dhiarmada agus Ghráinne ’The pursuit of Diarmaid and Grainne’ and the vast number of places claiming been touched by them (the several "beds" (Leaba Dhiarmada agus Ghráinne) across the country for example) on their flight^{5} shows the desire of the people to be associated with such a powerful character.

Clonmacnoise (Cluain Maccu Nois) ’Meadow of the sons of Nos’. This Nos was the son of Fiodhach, a chief of the tribe of Dealbhna, in whose territory Clonmacnoise was situated.

Clonshee (Cluain Sidh) ’Meadow of the Fairies’. As it is said that the places of fairies, their meadows, trees, caves, hills, etc. have to stay untouched and are not to be entered, this may have been a name for an especially beautiful and untouched area.

Cooley Hills (Cuailgne). After the defeat of the Tuatha De Dananns by the Milesians at Teltown in Meath (see below), the Milesian chief Cuailgn was slain here as he pursued the fleeing.

Loch Derg (Lough Dergdherc) ’lake of the red eye’ as it was called in reference to an old legend.

Seefin (Suidhe Finn) ’Seat of Finn Mac Coole’.

Teltown (Tailtenn) ’(place) of Taillte’. King Lui (Lewy) Of The Long Hand, one of the Tuatha De Danann, established a fair or gathering (now known as the festival of Lunasa) to be held here yearly at the first of August. It is a marry feast with games, pastimes and marriages celebrated. In honour of his foster mother Taillte this place came to its name.

Tara (Teamhair) ’residence on elevated spot’. Many places of this name exist in Ireland besides the famous Tara in Meath.

Templeshanbo (Seán both Sine) ’Temple (at the place) of Sin’s old hut’. The temple was only erected later in Christian times after the hut of Sin was long gone.

5See [2].

Bibliography

[1] Gerler Oliver, Lecture notes to GA4116, University of Limerick, 2000

[2] Gerler Oliver, A Discussion of the magical attributes of the Hero in Fenian Literature with specific reference to the tale ’The Pursuit of Diarmaid and Gráinne’, University of Limerick, 1999

[3] Joyce P.W., Pocket Guide to Irish Place Names, appletree, 1984, ISBN 0-86281-127-9

[4] Joyce P.W., Irish Names of Places, Scolar Press, vol.1, 4th ed., 1875, repr. 1972, ISBN 0-85409-916-6; vol.2, 1973, ISBN 0-85409-874-7; vol.3, 1976, ISBN 0-7158-1157-6

[5] Field John, Place Names of Great Britain and Ireland, Barnes & Noble, 1980, ISBN 0-389-20154-5

[6] An Coimisiún Logainmneacha, Sráidainmneacha: Treoirlínte, 1992